OpenStudy (anonymous):

Factor completely. -(x-y)^2 / (y^2-x^2)

rvc (rvc):

(a-b)^2=?

OpenStudy (anonymous):

im sorry I don't understand.

rvc (rvc):

\[(a-b)^2=a^2-2ab+b^2\] correct?

OpenStudy (anonymous):

oh I see.... yes it does.

rvc (rvc):

@chris117 got the ans?

OpenStudy (anonymous):

\[\frac{ 2xy-y^2 }{ y^2 } ?\]

rvc (rvc):

@hartnn pls help him im leaving

hartnn (hartnn):

lets take the denominator , apply the formula \(A^2-B^2 = (A+B)(A-B)\) so what will be y^2-x^2 = ... ?

OpenStudy (anonymous):

\[(x+y)(x-y)\]

hartnn (hartnn):

thats correct, and we write \(C^2 = C\times C \) so, \((x-y)^2 = (x-y)(x-y)\) right ?? put them together, anything getting cancelled from numerator and denominator ?

OpenStudy (anonymous):

from the original numerator the \[(x-y)\] ?

hartnn (hartnn):

oh and actually, \(y^2 -x^2 =(y+x)(y-x) \)

OpenStudy (anonymous):

@hartnn you asked chris for y^2-x^2 right? @chris117 said it is (x+y)(x-y) but isnt it (y+x)(y-x) ?

hartnn (hartnn):